OpenStudy (anonymous):

The volume of two similar solids is 1331 m3 and 729 m3. The surface area of the larger solid is 605 m2. What is the surface area, in square meters, of the smaller solid? 121 305 405 81

OpenStudy (callisto):

Since they are similar \[\sqrt{\frac{A_1}{A_2}} =\sqrt[3]{\frac{V_1}{V_2}}\]\[\sqrt{\frac{605}{A}} =\sqrt[3]{\frac{1331}{729}}\]\[\sqrt{\frac{605}{A}} =\frac{11}{9}\] Can you get A ( surface area of the smaller solid ) from here?

OpenStudy (anonymous):

I still do not fully understand , I think the answer is 405

OpenStudy (callisto):

The answer is 405.. For similar figure, (ratio of the length)^2 = ratio of the area (ratio of the length)^3 = ratio of the volume So, take square of area and take cubic root of volume, you'll get the ratio of the sides, which should be equal.